40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 25 15:41:29 2017, Lydia, Update, IMC, 29.5 MHz modulation depth measurement plan 
    Reply  Wed Jan 25 17:30:03 2017, Koji, Update, IMC, 29.5 MHz modulation depth measurement plan 
       Reply  Wed Jan 25 19:39:07 2017, gautam , Update, IMC, 29.5 MHz modulation depth measurement plan 
          Reply  Fri Jan 27 21:25:11 2017, Lydia, Update, IMC, 29.5 MHz modulation depth  
             Reply  Sat Jan 28 01:25:51 2017, gautam, Update, IMC, 29.5 MHz modulation depth  IMC_mod.pdfLockMCNew.png
                Reply  Mon Jan 30 19:07:48 2017, gautam, Update, IMC, RF AM stabilization box pulled out 
                   Reply  Tue Jan 31 01:07:20 2017, Lydia, Update, IMC, RF AM stabilization box pulled out AM_stablilizer_annotation.pdf
                      Reply  Tue Jan 31 13:46:34 2017, rana, Update, IMC, RF AM stabilization box pulled out 
                         Reply  Tue Jan 31 14:17:48 2017, gautam, Update, IMC, RF AM stabilization box pulled out 
                            Reply  Tue Jan 31 22:07:13 2017, gautam, Update, IMC, RF AM stabilization box revamp 
                            Reply  Tue Jan 31 22:28:39 2017, Lydia, Update, IMC, RF AM stabilization box pulled out AM_noise.pdf
                               Reply  Wed Feb 1 11:51:19 2017, Koji, Update, IMC, RF AM stabilization box pulled out 
                                  Reply  Wed Feb 1 16:45:56 2017, Lydia, Update, IMC, RF AM stabilizer box Modification Plan RF_AM_stabilizer_modification.pdf
                                     Reply  Wed Feb 1 20:49:34 2017, rana, Update, IMC, RF AM stabilizer box Modification Plan 
                                        Reply  Wed Feb 1 23:13:30 2017, Lydia, Update, IMC, RF AM stabilizer box Modification Plan 
                                           Reply  Sun Feb 5 21:56:50 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement 
                                              Reply  Tue Feb 7 12:01:10 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement RF_amp.pdf
                                                 Reply  Tue Feb 7 17:00:55 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement 
                                                    Reply  Fri Feb 10 02:14:10 2017, gautam, Update, IMC, 29.5 MHz stabilizer box replacement 
                                                       Reply  Fri Feb 10 11:41:43 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement 
                                                          Reply  Fri Feb 10 13:24:28 2017, rana, Update, IMC, 29.5 MHz stabilizer box replacement 
                                                             Reply  Fri Feb 10 18:21:21 2017, gautam, Update, IMC, IMC Demod board 
                                                                Reply  Fri Feb 10 19:32:15 2017, Koji, Update, IMC, IMC Demod board 
                                                                   Reply  Sun Feb 12 01:16:57 2017, gautam, Update, IMC, IMC length loop - summary of changes DemodBoardOrthogonality.pdfIMC_PDH.pdfTEK00001.PNGIMC_OLTF.pdfFSS_gain_comparison.pdf
                                                                      Reply  Mon Feb 13 11:55:14 2017, rana, Update, IMC, IMC length loop - summary of changes 
                                                                      Reply  Mon Feb 13 13:34:44 2017, gautam, Update, IMC, IMC length loop - bad SMA cable replaced noisy_PMC_REFL.png
                                                                         Reply  Mon Feb 13 19:44:55 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box rfAmp.pdf
                                                                            Reply  Wed Mar 1 21:15:40 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box RF_amp_(1).pdf
                                                                               Reply  Wed Mar 1 23:56:09 2017, gautam, Update, IMC, Front panel for 29.5 MHz amplifier box 
                                                                                  Reply  Thu Mar 2 20:32:18 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box 
                                                                         Reply  Wed Feb 15 23:54:13 2017, gautam, Update, IMC, IMC saga continues... MCstrip.pngIMCweirdness.png
                                                                            Reply  Fri Feb 17 20:10:18 2017, gautam, Update, IMC, WFS servos turned back on 
                                                                               Reply  Sat Feb 18 14:09:06 2017, rana, Update, IMC, WFS servos turned back on 
                                                                                  Reply  Sat Feb 18 21:50:48 2017, gautam, Update, IMC, WFS servos turned back on WFS_error_noise.pdf
                                                                                     Reply  Sun Mar 5 12:41:23 2017, gautam, Update, IMC, WFS servo-steppin WFS_stepping.pngWFS_comparisons.pdfWFSdigitalAA.pdf
Message ID: 12833     Entry time: Wed Feb 15 23:54:13 2017     In reply to: 12824     Reply to this: 12838
Author: gautam 
Type: Update 
Category: IMC 
Subject: IMC saga continues... 

Following the discussion at the meeting today, I wanted to finish up the WFS tuning and then hand over the IFO to Johannes for his loss stuff. So I did the following:

  1. First I set the dark offsets on the WFS (with PSL shutter closed). Then I hand aligned the MC to maximize transmission, centered the beam on the WFS, and set the RF offsets with the MC unlocked.
  2. Given that the demod phase for the IMC PDH demodulation board changed by |45 degrees|, I tried changing the digital demod phases in each of the WFS quadrant signals by +/- 45 degrees. Turns out +45 degrees put all the error signal into the I Phase, which is what we use for the WFS loops.
  3. Then I attempted to check the WFS loops. I estimated that we have ~25 times the modulation depth now, so I reduced the WFS1/2 P/Y gains by this factor (but left the MC2 TRANS P/Y gains as is). The loop gain seemed overall too low, so I upped the gain till I saw instability in the loop (error signals ringing up). Then I set the loop gains to 1/3 of this value - it was 0.01 before, and I found the loop behaved well (no oscillations, MC TRANS stabilized) at a gain of 0.002.

At this point, I figured I would leave the WFS in this state and observe its behaviour overnight. But abruptly, the IMC behaviour changed dramatically. I saw first that the IMC had trouble re-acquiring lock. Moreover, the PC Drive seemed saturated at 10.0V, even when there was no error signal to the MC Servo board. Looking at the MEDM screen, I noticed that the "C1-IOO_MC_SUM_MON" channel had picked up a large (~3V) DC offset, even with In1 and In2 disabled. Moreover, this phenomenon seemed completely correlated with opening/closing the PSL shutter. Johannes and I did some debugging to make sure that this wasn't a sticky button/slider issue, by disconnecting all the cables from the front panel of the servo board - but the behaviour persisted, there seemed to be some integration of the above-mentioned channel as soon as I opened the PSL shutter.

  

 

Next, I blocked first the MC REFL PD, and then each of the WFS - turns out, if the light to WFS2 was blocked and the PSL shutter opened, there was no integrating behaviour. But still, locking the MC was impossible. So I suspected that something was wrong with the LO inputs to the WFS Demod Boards. Sure enough, when I disconnected and terminated those outputs of the RF distribution box, I was able to re-lock the MC fine.

I can't explain this bizzare behaviour - why should an internal monitor channel of the MC Servo board integrate anything when the only input to it is the backplane connector (all front panel inputs physically disconnected, In1 and In2 MEDM switches off)? Also, I am not sure how my work on the WFS could have affected any hardware - I did not mess around at the 1X1 rack in the evening, and the light has been incident on the WFS heads for the past few days. The change in modulation depth shouldn't have resulted in the RF power in this chain crossing any sort of damage threshold since the measured power before the changes was at the level of -70dBm, and so should be at most -40dBm now (at the WFS demod board input). The only thing different today was that the digital inputs of the WFS servos were turned on...

So for tonight I am leaving the two outputs of the RF distribution box that serve as the LO for the WFS demod boards terminated, and have also blocked the light to both WFS with beam blocks. The IMC seems to be holding lock steady, PC drive levels look normal...


Unrelated to this work, but I have committed to the svn the updated versions of the mcup and mcdown scripts, to reflect the new gains for the autolocker...

ELOG V3.1.3-